The equation of the line that goes through the point ( 5, 5 ) and is parallel to the line going through the points ( -4 ,6 ) and ( 2 ,4 ) can be written in the form y = mx+b where m is and b is. I am struggling with this question i have no idea how to start this can you explain it in steps for me.
Answer by ewatrrr(10682)

 
Hi,
The equation of the line that goes through the point ( 5, 5 )
and is parallel to the line going through the points
( -4 ,6 ) and
( 2 ,4 ) m = 2/-6 = (-1/3)
New Line: (same slope)
y = mx + b
y = (-1/3)x + b ||using Pt(5,5) to solve for b
5 = -5/3 + b
20/3 = b
y = (-1/3)x + 20/3